OpenStudy (anonymous):

A regular polygon has 15 sides. Find the measure of each angle.

OpenStudy (anonymous):

Interior or exterior angle?

OpenStudy (anonymous):

It never said.

OpenStudy (anonymous):

I would assume they want interior. A square is a regular polygon, 4 sides, interior and exterior angles are 90. A pentagon is a regular polygon, 5 sides. Exterior angle = 360/5=72 Interior angle = 180-72 = 108 Hexagon...6 sides... exterior angle is 360/6 = 60, interior is 180-60 = 120 Continue for 15 sides... exterior angle = 360/15, interior = 180-exterior.